A plane travels from city A to city B and it takes two hours...on the second trip the plane travels from city A to City B and it flys 10mph slower and it takes 2 1/2 hours...what is the distance from city A to city B?
Answer by ankor@dixie-net.com(12704)

A plane travels from city A to city B and it takes two hours...on the second trip the plane travels from city A to City B and it flies 10 mph slower and it takes 2 1/2 hours...what is the distance from city A to city B?
:
The says essentially two trips of the same distance at two different speeds
:
Let s = speed on the 2 hr trip
Then (s-10) = speed on the 2.5 hr trip
:
Write a distance equation: Distance = time * speed
:
First trip dist = 2nd trip dist
2s = 2.5(s-10)
2s = 2.5s - 25
2s - 2.5s = - 25
-.5s = -25
s = -25/-.5
s = + 50 mph
:
Find the distance using 50 mph for 2 hrs
50 * 2 = 100 mi
:
Check the solution using 40 mph for 2.5 hrs
40 * 2.5 = 100 m